### v1

# HiVG: Hierarchical Multimodal Fine-grained Modulation for Visual Grounding

Abstract.
Visual grounding, which aims to ground a visual region via natural language, is a task that heavily relies on cross-modal alignment. Existing works utilized uni-modal pre-trained models to transfer visual/linguistic knowledge separately while ignoring the multimodal corresponding information. Motivated by recent advancements in contrastive language-image pre-training and low-rank adaptation (LoRA) methods, we aim to solve the grounding task based on multimodal pre-training. However, there exists significant task gaps between pre-training and grounding. Therefore, to address these gaps, we propose a concise and efficient hierarchical multimodal fine-grained modulation framework, namely HiVG. Specifically, HiVG consists of a multi-layer adaptive cross-modal bridge and a hierarchical multimodal low-rank adaptation (Hi LoRA) paradigm. The cross-modal bridge can address the inconsistency between visual features and those required for grounding, and establish a connection between multi-level visual and text features. Hi LoRA prevents the accumulation of perceptual errors by adapting the cross-modal features from shallow to deep layers in a hierarchical manner. Experimental results on five datasets demonstrate the effectiveness of our approach and showcase the significant grounding capabilities as well as promising energy efficiency advantages. The project page: https://github.com/linhuixiao/HiVG.
Multimodality; Visual Grounding; Referring Expression Comprehension; Low-Rank Adaptation; Hierarchical
copyright: acmlicensedjournalyear: 2024conference: Proceedings of the ACM International Conference on Multimedia; 2024; ArXivisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Visual Grounding (VG), also known as Referring Expression Comprehension (REC) or Phrase Grounding (Qiao et al., 2020; Mao et al., 2016; Yu et al., 2016; Hu et al., 2016; Deng et al., 2021; Xiao et al., 2023, 2019; Li et al., 2019), is a fundamental and challenging task at the intersection fields of vision-language understanding, which can be potentially used in a wide range of applications, such as visual question answering (Antol et al., 2015) and human-machine interaction (Chen et al., 2023b) etc.. Unlike object detection (Liu et al., 2023b, c), which requires a predefined and fixed set of categories, grounding is not limited to specific categories but instead needs to identify the specific image region according to the language expression semantics. Thus, grounding is a task that strongly relies on the interaction and alignment of multimodal features.
Existing state-of-the-art (SOTA) approaches (Deng et al., 2021; Ye et al., 2022a; Deng et al., 2023; Su et al., 2023; Ho et al., 2023; Zhao et al., 2022; Zhu et al., 2022) utilize uni-modal pre-trained detection models or language models (e.g., ResNet (He et al., 2016), Swin Transformer (Liu et al., 2021), DETR (Carion et al., 2020), ViT-Det (Li et al., 2022a), BERT (Devlin et al., 2018), RoBERTa (Liu et al., 2019a) etc.) to facilitate grounding learning. These methods separately transfer the language or vision knowledge from pre-trained models by using resource-consuming fully parameter fine-tuning, ignoring the multimodal corresponding information. Therefore, it is natural for us to consider using cross-modal pre-trained models as a solution to the grounding problem.
Figure 1. Visual attentions and grounding results of CLIP and the proposed HiVG. The attentions are perceived by the [CLS] token over vision tokens.
By utilizing language supervision from large-scale unlabeled data, Vision-Language Pre-training (VLP) can acquire comprehensive multimodal representations. Recently, the remarkable success of Contrastive Language-Image Pre-training (CLIP) (Radford et al., 2021) has demonstrated its ability to learn general visual concepts, which assists many multimodal tasks to achieve remarkable improvements (Radford et al., 2021; Peng et al., 2023b; Tao et al., 2023; Kim et al., 2023; Wang et al., 2022a). In visual grounding, there are also works, e.g., CLIP-VG (Xiao et al., 2023) and Dynamic-MDETR (Shi et al., 2022), which consider using CLIP. However, existing methods mainly utilize the CLIP as a backbone to extract strong vision and language features, without comprehensively investigating on the significant task gap between the pre-trained CLIP and the downstream grounding, which hinders exploiting the full potential of pre-training models. In this work, we scrutinize the task gap from two aspects. (1) Data bias. There inevitably exists a certain bias in data between the large-scale pre-training and grounding. Directly utilizing the frozen vision backbone of the CLIP may extract visual features sensitive to general objects that are not the focus of the query in visual grounding. For example, as shown in Fig. 1-(a), the middle giraffe receives highlight attention, but it has little relation to the grounding task. (2) Difference in learning objectives. The visual grounding task needs to find the precise image region that has the target object expressed by the query sentence. In contrast, CLIP works as a multimodal pre-trained model, which is only constrained to coarsely align noisy image and text data (Schuhmann et al., 2021) in a self-supervised way. In addition, the self-supervised constraint is only performed at the final layer. When directly using the pre-trained CLIP in visual grounding, some valuable fine-grained visual information in the bottom vision layers may be discarded, which brings challenges for accurately locating the object box. For example, as shown in Fig. 1-(a), the left giraffe receives relatively small attention areas, which leads to inaccurate box of the target object.
It is not trivial to address the two kinds of task gaps. (1) For the task gap of data bias, extracting the features of the query text to guide the visual feature learning is a potential way to solve it. However, the query text has a feature space that is very different from the visual space and it is difficult to find the appropriate semantic information from the query features to guide the learning of different vision layers. (2) For the task gap of learning objectives, to adapt the pre-trained CLIP to the grounding task, a straightforward way is fine-tuning the pre-trained weights. Whereas, this scheme may lead to catastrophic forgetting, which is harmful to retain the general knowledge learned by the pre-trained models. Another potential solution is to employ Low-Rank Adaptation (LoRA) (Hu et al., 2021) by fine-tuning only a few parameters. However, simply applying LoRA does not achieve fine-grained adaptation and even lead to performance degradation. Since high-level features depend on low-level features, and they are susceptible to perturbations of the shallow features. If all layers of a large-scale pre-trained model are adapted simultaneously, perceptual errors in bottom layers may accumulate and amplify. Therefore, it is necessary to consider a hierarchical approach for progressively adapt fine-grained visual features from shallow to deep layers.
In this paper, we propose a hierarchical multimodal fine-grained modulation framework to more effectively adapt the pre-trained CLIP to grounding, namely HiVG. It is a concise and efficient end-to-end framework that can simultaneously alleviate two kinds of task gaps (i.e., data bias and learning objectives) through a multi-layer adaptive cross-modal bridge and a hierarchical low-rank adaptation paradigm.
Firstly, to address the inconsistency between visual features of the pre-trained CLIP and those required for grounding, as well as establish a connection between multi-level visual and text features, we have designed a multi-layer adaptive cross-modal bridge. Specifically, the cross-modal bridge includes a sample-agnostic semantic weighting module and a multi-head cross-attention module. The weighting module incorporates learnable multi-level sample-agnostic adaptive weights, facilitating the selection of appropriate linguistic features through a residual operation. The multi-head cross-attention utilizes the selected multi-level text features for guiding the learning of the visual features required in grounding. The sample-agnostic semantic weighting scheme is inspired by (Cao et al., 2020), i.e., specific layers of a pre-trained model may have distinct responses to certain concepts or semantics that are independent of the input and relevant to the network layers.
Secondly, to prevent the accumulation of errors layer by layer in the downstream adaptation process of the pre-trained model, we propose Hierarchical Low-rank Adaptation (Hi LoRA) paradigm. Existing methods mainly utilize LoRA(Hu et al., 2021) as a parameter-efficient fine-tuning (PEFT) method to learn a single round along with the entire model. Different from previous methods (Hu et al., 2021; Smith et al., 2023), we divide the network layers of the pre-trained CLIP into multiple layer groups. The low-rank adaptation is allocated into multiple stages where each stage relates to several layer groups. Then, during the adaptation process, visual features are recursively and hierarchically adapted from shallow to deep layers in a hierarchical manner. Simultaneously, with the assist of the multi-layer cross-modal bridge, Hi LoRA can not only achieve fine-grained hierarchical adaptation, but also enable the low-rank matrix perception based on the vision and language cross-modal information.
[…]
The main contributions can be summarized as three-fold:
  • We proposed a concise hierarchical multimodal modulation framework, which utilizes the hierarchical structure to gradually adapt CLIP to grounding. HiVG achieves fine-grained interaction between multi-level visual representations and language semantics, and significantly alleviates the task gap between CLIP and grounding.
  • We are the first to propose the hierarchical multimodal low-rank adaptation structure. Hi LoRA is a basic and concise hierarchical adaptation paradigm, which is task-agnostic.
  • We conducted extensive experiments to verify the effectiveness of HiVG approaches. Results show that our method achieves promising results, surpassing the SOTA methods under the same setting by a significant margin. Besides, our model offers significant computing efficiency advantages.
2. Related Work
2.1. Visual Grounding
Visual grounding has recently received significant research attention, and it can be categorized into several settings. On the one hand, represented by TransVG (Deng et al., 2021), this setting involves full-parameter fine-tuning utilizing pre-trained closed-set detectors and language models. It is considered the most conventional and extensively studied setting. Under this setting, numerous complex two-stage (Yu et al., 2018; Liu et al., 2019b, c; Hong et al., 2019) and one-stage (Yang et al., 2020, 2019a; Zhou et al., 2021) methods emerged based on traditional detection networks in the early CNN era. After the introduction of ViT (Vaswani et al., 2017; Dosovitskiy et al., 2020), the Transformer-based networks (Deng et al., 2021; Ye et al., 2022a; Kamath et al., 2021; Ho et al., 2023; Li and Sigal, 2021; Xiong et al., 2023; Su et al., 2023; Deng et al., 2023; Ye et al., 2022b) constantly pushes the accuracy to new limits. However, these works only focus on achieving grounding by using independently pre-trained uni-modal detectors and language encoders while ignoring the alignment of cross-modality information within pre-trained model itself. More recent works, such as QRNet (Ye et al., 2022a), VG-LAW (Su et al., 2023), TransVG++ (Deng et al., 2023), etc., only incorporate language-guided knowledge in vision backbone without attempting multi-level fine-grained alignment of multimodal features. Motivated by this setting, several works, such as CLIP-VG (Xiao et al., 2023) and Dynamic-MDETR (Shi et al., 2022), have recently sprung up to the setting of fine-tuning with vision and language self-supervised pre-trained models. Following this setting, our study delves into a deeper perspective of hierarchical multimodal information and achieves fine-grained interaction of cross-modal features. On the other hand, with the evolution of the pre-training paradigm, many new settings have recently emerged that significantly improve the grounding performance, such as fine-tuning with box-level dataset-mixed open-set detection pre-trained models (e.g., MDETR (Kamath et al., 2021), Grounding-DINO (Liu et al., 2023d), etc.), fine-tuning with box-level / multi-task mixup-supervised pre-trained models (e.g., UniTAB (Yang et al., 2022), UNITER (Chen et al., 2020), OFA (Wang et al., 2022b), ONE-PEACE(Wang et al., 2023b), etc.), and grounding multimodal large language models (GMLLM, e.g., Shikra (Chen et al., 2023b), Kosmos-2 (Peng et al., 2023a), Ferret (You et al., 2024), LION (Chen et al., 2023a), etc.). However, these works require a large amount of fine-grained labeled data, resulting in a relatively high training cost.
2.2. Contrastive Language-Image Pre-training
With the promotion of learning general and transferable cross-modal representations, VLP has become the core training paradigm and research foundation of modern Vision-Language (V-L) research. Benefiting from self-supervised contrastive learning, CLIP has demonstrated impressive generalization and downstream transfer ability in a series of studies (Peng et al., 2023b; Tao et al., 2023). More recently, some works utilized CLIP to realize grounding transfer, such as adapting-CLIP (Li et al., 2022b), ReCLIP (Subramanian et al., 2022) etc., but these works are limited to using CLIP features as aids in an unsupervised or zero-shot setting and cannot directly perform grounding. Although CLIP-VG (Xiao et al., 2023), Dynamic-MDETR (Shi et al., 2022), JMR (Zhu et al., 2023) etc., realizes grounding transfer, it does not conduct more in-depth research on the task gaps and the hierarchical cross-modal features. Unlike previous work, our study fills the gap by conducting a more comprehensive study of the cross-modal task gaps between CLIP’s pre-training and downstream grounding.
2.3. Low-Rank Adaptation
LoRA (Hu et al., 2021) freezes the pre-trained model weights and injects trainable rank decomposition matrices into each layer of the Transformer architecture, thereby significantly reducing the number of trainable parameters for downstream tasks. Vanilla LoRA has been proposed in the field of natural language processing for Large Language Models (LLM) such as LLaMA2 (Touvron et al., 2023), GPT-2 (Radford et al., 2019), GPT-3 (Brown et al., 2020) with 175B parameters, etc.. Recently, researchers have attempted to apply vanilla LoRA in the fields of cross-modal tasks (Smith et al., 2023). However, since cross-modal tasks primarily emphasize the interaction of multimodal information in contrast to unimodal language or visual tasks, the application of LoRA to grounding tasks remains unexplored. Consequently, we propose Hi LoRA as an effective solution for addressing the existing gaps in multimodal downstream transfer.
3. Methodology
Figure 2. Schematic representation of the hierarchical multimodal fine-grained modulation framework.
In this section, we propose our hierarchical multimodal fine-grained modulation framework for visual grounding, namely HiVG, which mainly consists of the multi-layer adaptive cross-modal bridge and the hierarchical low-rank adaptation (Hi LoRA) paradigm. We will introduce each of these methods in the following sections.
3.1. Framework Overview
Our aim is to achieve fine-grained hierarchical cross-modal feature modulation, so as to narrow the task gap between the self-supervised pre-training and grounding. Therefore, we integrate the multi-level image and text representations from a hierarchical perspective with the facilitation of multi-layer adaptive cross-modal bridge and the hierarchical LoRA paradigm. Specifically, as shown in Fig. 2, the network architecture of HiVG consists of a CLIP image encoder, a CLIP text encoder, a grounding encoder and a regression head. Firstly, for any given image $\mathcal{I}\in\mathbb{R}^{3\times H\times W}$ and text $\mathcal{T}\in\mathbb{R}^{L_{l}}$ pairs, the visual and text encoders encode the image and text tokens to obtain the visual feature $\bm{f}_{v}\in\mathbb{R}^{L_{v}\times H_{v}}$ and text feature $\bm{f}_{l}\in\mathbb{R}^{L_{l}\times H_{l}}$ , respectively, where $H,W$ are the image size, $H_{v}$ and $H_{l}$ are the visual and text hidden embedding dimension, $L_{v}$ is the length of image token, which tokenized by a convolution projection, and $L_{l}$ is the length of text token, which tokenized by a lower-cased Byte Pair Encoding (BPE) with a 49,152 vocab size (Sennrich et al., 2016). We extract the multi-level intermediate visual features $\{\bm{f}_{v}^{i}\}_{i=1}^{m}\in\mathbb{R}^{m\times L_{v}\times H_{v}}$ and text features $\{\bm{f}_{l}^{i}\}_{i=1}^{n}\in\mathbb{R}^{n\times L_{l}\times H_{l}}$, which obtained by the ViT block and text Transformer block, respectively, where $m$ and $n$ are the numbers of extracted layers.
Simultaneously, to reduce the inconsistency between the visual features of the uni-modal image backbone and those required for grounding, we introduce a multi-layer adaptive cross-modal bridge to the visual encoder that bridges image and text modalities. Each layer of the bridge has a learnable sample-agnostic weighting module, thus enabling the uni-modal visual backbone to perceive hierarchical cross-modal text features.
Additionally, to prevent the accumulation and amplification of perceptual errors in the visual encoder, we propose a hierarchical low-rank adaptation (Hi LoRA) paradigm to adapting the pre-trained frozen parameters. During Hi LoRA training, the entire adaptation process learns from shallow to deep layers. The gradients backward from the grounding encoder are updated hierarchically and adaptively into the low-rank matrix based on both visual features and hierarchical language features. Besides, the intermediate visual features are aggregated and fed to the grounding encoder, which not only benefits the perception of multi-level visual features but also facilitates direct gradient backward updates without going from deep to shallow in the Hi LoRA low-stage training.
Finally, in the grounding encoder, we concatenate the multi-level visual features along with the hidden dimension, and leverage the weight $W_{mvp}\in\mathbb{R}^{(m\cdot H_{v})\times H_{g}}$ of a MLP-based visual perceiver to project them into embedding space $\bm{g}_{v}\in\mathbb{R}^{L_{v}\times H_{g}}$ with dimension $H_{g}$ to perceive multi-level visual representations:
(1) ${\bm{g}_{v}={\rm{concat}}[\bm{f}_{v}^{1},\bm{f}_{v}^{2},\cdots,\bm{f}_{v}^{m}]\otimes W_{mvp}}.\vspace{-2pt}$
To prevent any perturbation on $[EOS]$ token and ensure the subsequent constraints remain unaffected, we exclusively utilize the linear projection features $\bm{g}_{l}\in\mathbb{R}^{L_{l}\times H_{g}}$ of the last layer’s text features $\bm{f}_{l}^{last}$ to feed the grounding encoder. Finally, the input tokens of the grounding encoder are as follows:
(2) $\small\bm{x_{g}}=[\mathscr{g}_{r},\ cls,\ \underbrace{g_{v}^{1},\ g_{v}^{2},\ g_{v}^{3},\ \cdots,\ g_{v}^{L_{v}}}_{\text{CLIP image tokens}~{}\bm{g}_{v}}\ ,\ \underbrace{g_{l}^{1},\ g_{l}^{2},\ g_{l}^{3},\ \cdots,\ g_{l}^{L_{l}}}_{\text{CLIP text tokens}~{}\bm{g}_{l}}],\vspace{-4pt}$
where $cls$ represents the classification token $[CLS]$, $g_{r}$ represents the learnable $[REG]$ token, which is used to output the regression results (Deng et al., 2021). The $[EOS]$ token is the end token of each sequence within $\bm{f}_{l}$ and $\bm{g}_{l}$. The regression head is employed to conduct bounding box regression, which is a three-layer MLPs (Deng et al., 2021), each consisting of a linear layer and a ReLU activation layer. It outputs the final coordinate of the predicted grounding box ^$\hat{\mathcal{B}}=(\hat{x},\hat{y},\hat{w},\hat{h})$.
3.2. Multi-layer Adaptive Cross-modal Bridge
The visual encoder of CLIP independently encodes the image, and the obtained multi-level visual features may be inconsistent with those required for grounding. Additionally, as inspired by (Cao et al., 2020), specific layers of a pre-trained model may exhibit distinct responses to certain concepts or semantics that are independent of the input and relevant to the network layers. Therefore, we should provide a wide range of multi-level text features for different visual layers to select and calibrate. Thus, to address these issues, we propose integrating a multi-layer adaptive cross-modal bridge into the image encoder to achieve fine-grained visual features.
The multi-layer adaptive cross-modal bridge (MACB) mainly consists of a sample-agnostic semantic weighting module and a multi-head cross-attention. It is inserted into specific ViT blocks, and we define the layer index set $C$ as the insertion positions. The sample-agnostic weighting module enables distinct hierarchical language feature perception among different layers. Specifically, we first extract and aggregate the intermediate language features $\{\bm{f}_{l}^{i}\}_{i=1}^{n}\in\mathbb{R}^{n\times L_{l}\times H_{l}}$. Then, to stably strengthen or weaken the text features preferred by different visual layers, we utilize a residual operation to achieve selection of multi-level text features:
(3) ${{}^{*}\bm{f}}_{l}^{i}=\bm{w}_{l}^{i}\odot\bm{f}_{l}^{i}+\bm{f}_{l}^{i}.\vspace{-2pt}$
where $\{\bm{w}_{l}^{i}\}_{i=1}^{n}\in\mathbb{R}^{n\times L_{l}\times H_{l}}$ represent the learnable multi-level sample-agnostic adaptive weights within different layers, which can promote different visual layers respond distinctly to specific textual concepts or semantics. The weighted features are obtained by dot product between the sample-agnostic adaptive weights and multi-level features. We then add the weighted features to the original features to obtain the calibrated text features $\{^{*}\bm{f}_{l}^{i}\}_{i=1}^{n}$. Subsequently, we concatenate and project them into visual embedding space ${{}^{*}\bm{f}}_{l}^{ml}\in\mathbb{R}^{L_{l}\times H_{v}}$ to perceive multi-level language representations with linear projection weight $W_{proj}\in\mathbb{R}^{(n\cdot H_{l})\times H_{v}}$:
(4) ${{{}^{*}\bm{f}}_{l}^{ml}={\rm{concat}}[{{}^{*}\bm{f}}_{l}^{1},{{}^{*}\bm{f}}_{l}^{2},\cdots,{{}^{*}\bm{f}}_{l}^{n}]\otimes W_{proj}}.$
Finally, we perform a multi-head cross-attention on the calibrated multi-level text features ${{}^{*}\bm{f}}_{l}^{ml}$ (as key and value) and the layer-normalized visual features outputted by the self-attention in the ViT block (as query). Then, we add the resulting semantic-aware visual features ${\bm{f}}_{v}^{sa}$ back to the block as residuals after a FFN operation.
Figure 3. Hi LoRA and vanilla LoRA. (a) The vanilla LoRA learns the global low-rank matrix utilizing the entire set of pre-trained weights in a single round. (b) The proposed Hi LoRA employs a hierarchical approach to adapt the pre-trained model in a progressive manner, thereby finely reducing the task gap between pre-training and transfer tasks.
3.3. Hierarchical Low-Rank Adaptation
Although the cross-modal bridge enables the visual encoder to incorporate language information, its residual connection manner cannot adapt the frozen parameters of the pre-trained model. As a result, there is still a discrepancy between the visual features and those required for grounding, which may lead to cumulative and amplified perceptual errors layer by layer. LoRA (Hu et al., 2021) presents a potentially feasible solution. However, as clarified in the Sec. 1, the vanilla LoRA performs one-round learning also cannot address these issues. To avoid cumulative and amplified perceptual errors, we need to design a hierarchical adaptation paradigm.
Instead of directly training specific dense layers in a neural network, vanilla LoRA (Hu et al., 2021) indirectly optimizes the rank-decomposition matrices of the changes occurring in dense layers while keeping the pre-trained weights frozen. As depicted in Fig. 3-(a), based on the vanilla LoRA definition, we can substitute the weight updates for a pre-trained weight $W_{0}\in\mathbb{R}^{d\times k}$ with a low-rank decomposition $W_{0}+\Delta W=W_{0}+BA$, where $B\in\mathbb{R}^{d\times r},A\in\mathbb{R}^{r\times k}$, and $r\ll\min(d,k)$, i.e., the low rank $r$ is much smaller than the dimension $(d,k)$ of the original model. Throughout training, $W_{0}$ remains frozen, while $A$ and $B$ encompass trainable parameters. For hidden state $h=W_{0}x$, the forward procedure can be formulated as:
(5) $h=W_{0}x+\Delta Wx=W_{0}x+BAx.\vspace{-3pt}$
To consider the hierarchical scenario, we first define two concepts, i.e., layer group and LoRA stage. Layer group represents the divisions of the pre-trained network layers, while LoRA stage represents the execution of a small LoRA operation. By dividing the network layers of the pre-trained model into multiple layer groups, the learning of LoRA is divided into multiple stages where each stage relates to several layer groups. As depicted in Fig. 3-(b), from a hierarchical perspective, Hierarchical LoRA (Hi LoRA) structure enables downstream task adaptation progressively from the shallow to deep layer within the network with multiple LoRA stages.
Specifically, we define the total layers of the pre-trained network as $L$, and then divide it into $G$ groups, each containing $L/G$ layers. Then, we denote $W_{0}^{l}\in\mathbb{R}^{d\times k}$ as the pre-trained weights of $l^{th}$ layer block in the network, where $l\in[1,L]$. We utilize LoRA $j$ ($1\leq j\leq G$) to represent the $j^{th}$ adaptation stage. We denote the low-rank matrices of Hi LoRA at the $j^{th}$ stage of the $l^{th}$ layer block as $A_{j}^{l}$ and $B_{j}^{l}$, and $A_{j}$ contains ${\{A_{j}^{l}\}_{l=1}^{j\cdot L/G}}$, $B_{j}$ contains ${\{B_{j}^{l}\}_{l=1}^{j\cdot L/G}}$, then each LoRA stage $j$ will update the low-rank matrices of $A_{j}$ and $B_{j}$. We denote ${h_{j}^{l}}$ as the hidden state $h$ at Hi LoRA $j^{th}$ stage of the $l^{th}$ layer block. Then, the forward process of Hi LoRA in each hidden state ${h_{j}^{l}}$ ($j\in[1,G]$) can be formulated as:
(6) $h_{j}^{l}=\begin{cases}W_{0}^{l}x^{l},\ \ \ \ when\ l>j\cdot L/G,\\ W_{0}^{l}x^{l}+\sum_{k=\lceil l\cdot G/L\rceil}^{j}B_{k}^{l}A_{k}^{l}x^{l},\ \ when\ l\leq j\cdot L/G,\end{cases}\vspace{-2pt}$
[…]
With the assistance of the hierarchical mechanism, we can achieve better multimodal low-rank adaptation of multi-level visual features by utilizing textual semantic-aware visual features provided by the adaptive cross-modal bridge. Specifically, the layer groups of Hi LoRA are associated with the insertion positions $C$ of the bridge. When $l>j\cdot L/G$, the forward process of Hi LoRA in each hidden state ${h_{j}^{l}}$ can be formulated as:
(7) $h_{j}^{l}=\begin{cases}W_{0}^{l}{\bm{f}_{v}}^{l-1},\ \ when\ l\notin C,\\ W_{0}^{l}{({\bm{f}_{v}}^{l-1}+{\bm{f}}_{v}^{sa})},\ when\ l\in C.\end{cases}\vspace{-2pt}$
While in $l\leq j\cdot L/G$, the process can be formulated as:
(8) $\small h_{j}^{l}=\begin{cases}W_{0}^{l}{\bm{f}_{v}}^{l-1}+\sum_{k=\lceil l\cdot G/L\rceil}^{j}B_{k}^{l}A_{k}^{l}{\bm{f}_{v}}^{l-1},\ \ when\ l\notin C,\\ W_{0}^{l}{{({\bm{f}_{v}}^{l-1}+{\bm{f}}_{v}^{sa})}}+\sum_{k=\lceil l\cdot G/L\rceil}^{j}B_{k}^{l}A_{k}^{l}{{({\bm{f}_{v}}^{l-1}+{\bm{f}}_{v}^{sa})}},\ when\ l\in C.\\ \end{cases}\vspace{-2pt}$
During the backward process, the updates are gradually performed from $1^{st}$ to $G^{th}$ stage, and the learning rate can vary at different stages. Additionally, we use a random Gaussian initialization for $A$ and 00 for $B$, so $\Delta W=BA$ is 00 at the beginning of training. We then scale $\Delta Wx$ by $\frac{\alpha}{r}$, where $\alpha$ is a constant in $r$. To mitigate inference latency or parameter increase, we incorporate the low-rank matrix into the pre-trained weights after every training stage.
Hi LoRA provides a new interaction for refining latent representation, preventing direct gradient propagation of vanilla LoRA from deep to shallow layers. Simultaneously, through its hierarchical mechanism, it can avoid the accumulation of perceptual errors in the fine-tuning process, enabling fine-grained cross-modal interaction. Finally, it is worth noting that Hi LoRA represents a basic hierarchical adaptation paradigm that is task-agnostic.
Table 1. Comparison with latest SOTA methods on RefCOCO/+/g (Yu et al., 2016; Mao et al., 2016), ReferItGame (Kazemzadeh et al., 2014a) and Flickr30k Entities (Plummer et al., 2015) for grounding task. * represents utilizing ImageNet (Krizhevsky et al., 2012) pre-training. $\dagger$ indicates that all of the RefCOCO/+/g training data has been used during pre-training. RN101, DN53, Swin-S, and ViT-B are shorthand for the ResNet101, DarkNet53, Swin-Transformer Small, and ViT Base, respectively. The latest CLIP-based SOTA methods are shaded in gray. We highlight the best performance of the base model in the red colors and bold the best results for the large model.
Methods Venue Visual Language Multi- RefCOCO RefCOCO+ RefCOCOg ReferIt Flickr
Backbone Backbone task val testA testB val testA testB val test test test
Fine-tuning w. uni-modal pre-trained close-set detector and language model: (traditional setting)
TransVG (Deng et al., 2021) ICCV’21 RN101+DETR BERT-B 81.02 82.72 78.35 64.82 70.70 56.94 68.67 67.73 70.73 79.10
SeqTR (Zhu et al., 2022) ECCV’22 DN53 BiGRU 81.23 85.00 76.08 68.82 75.37 58.78 71.35 71.58 69.66 81.23
RefTR* (Li and Sigal, 2021) NeurIPS’21 RN101+DETR BERT-B 82.23 85.59 76.57 71.58 75.96 62.16 69.41 69.40 71.42 78.66
Word2Pix (Zhao et al., 2022) TNNLS’22 RN101+DETR BERT-B 81.20 84.39 78.12 69.74 76.11 61.24 70.81 71.34
YORO (Ho et al., 2023) ECCV’22 ViLT (Kim et al., 2021) BERT-B 82.90 85.60 77.40 73.50 78.60 64.90 73.40 74.30 71.90
QRNet (Ye et al., 2022a) CVPR’22 Swin-S(Liu et al., 2021) BERT-B 84.01 85.85 82.34 72.94 76.17 63.81 71.89 73.03 74.61 81.95
VG-LAW (Su et al., 2023) CVPR’23 ViT-Det (Li et al., 2022a) BERT-B 86.06 88.56 82.87 75.74 80.32 66.69 75.31 75.95 76.60
TransVG++(Deng et al., 2023) TPAMI’23 ViT-Det (Li et al., 2022a) BERT-B 86.28 88.37 80.97 75.39 80.45 66.28 76.18 76.30 74.70 81.49
Fine-tuning w. vision-language self-supervised pre-trained model:
CLIP-VG (Xiao et al., 2023) TMM’23 CLIP-B CLIP-B 84.29 87.76 78.43 69.55 77.33 57.62 73.18 72.54 70.89 81.99
JMRI (Zhu et al., 2023) TIM’23 CLIP-B CLIP-B 82.97 87.30 74.62 71.17 79.82 57.01 71.96 72.04 68.23 79.90
Dynamic-MDETR TPAMI’23 CLIP-B CLIP-B 85.97 88.82 80.12 74.83 81.70 63.44 74.14 74.49 70.37 81.89
HiVG (ours) CLIP-B CLIP-B 87.32 89.86 83.27 78.06 84.81 68.11 78.29 78.79 75.22 82.11
HiVG-L (ours) CLIP-L CLIP-L 88.14 91.09 83.71 80.10 86.77 70.53 80.78 80.25 76.23 82.16
Fine-tuning w. box-level dataset-mixed open-set detection pre-trained model / multi-task mix-supervised pre-trained model:
MDETR (Kamath et al., 2021) ICCV’21 RN101+DETR RoBERT-B 86.75 89.58 81.41 79.52 84.09 70.62 81.64 80.89 83.80
DQ-DETR (Liu et al., 2023a) AAAI’23 RN101+DETR BERT-B 88.63 91.04 83.51 81.66 86.15 73.21 82.76 83.44
Grounding-DINO Arxiv’23 Swin-T BERT-B 89.19 91.86 85.99 81.09 87.40 74.71 84.15 84.94
UniTAB (Yang et al., 2022) ECCV’22 RN101+DETR RoBERT-B 86.32 88.84 80.61 78.70 83.22 69.48 79.96 79.97 79.38
OFA-B (Wang et al., 2022b) ICML’22 OFA-B OFA-B 88.48 90.67 83.30 81.39 87.15 74.29 82.29 82.31
OFA-L (Wang et al., 2022b) ICML’22 OFA-L OFA-L 90.05 92.93 85.26 85.80 89.87 79.22 85.89 86.55
HiVG (ours) CLIP-B CLIP-B 90.56 92.55 87.23 83.08 89.21 76.68 86.52 86.62 77.75 82.08
HiVG-L (ours) CLIP-L CLIP-L 91.37 93.64 88.03 86.78 89.91 78.02 88.61 88.60 78.16 82.63
3.4. Training Objectives
To ensure the features learned by the cross-modal hierarchical structure meet the fine-grained and regional properties, we design multiple constraints to facilitate the training of HiVG framework.
Contrastive Learning Constraint. To enhance training stability, we employ image-text Contrastive Learning (CL) as a constraint for Hi LoRA. CL can also be formed between the grounding expression and the images within a shuffled training batch when differences are adequate. We treat the grounding image-text pairs as positive and all other random pairs as negative. We minimize the sum of two losses, one for text-to-image matching:
(9) $\mathcal{L}_{t2i}=-\frac{1}{N}\sum_{i}^{N}\log{\frac{\exp(<\bm{t}_{i}^{\top},\bm{v}_{i}>/\tau)}{\sum_{j=1}^{N}\exp(<\bm{t}_{i}^{\top},\bm{v}_{j}>/\tau)}},\vspace{-3pt}$
and the other for image-to-text matching:
(10) $\mathcal{L}_{i2t}=-\frac{1}{N}\sum_{i}^{N}\log{\frac{\exp(<\bm{v}_{i}^{\top},\bm{t}_{i}>/\tau)}{\sum_{j=1}^{N}\exp(<\bm{v}_{i}^{\top},\bm{t}_{j}>/\tau)}},\vspace{-3pt}$
where $N$ is the batch size, $\bm{v}_{i}$ and $\bm{t}_{j}$ are the normalized embeddings of image in $i^{th}$ pair and that of text in $j^{th}$ pair, respectively. $\tau$ is the temperature to scale the logits, and <,>$\cdot$,$\cdot$>< ⋅ , ⋅ > denotes cosine similarity operation. Therefore, the constraint can be formulated as:
(11) $\mathcal{L}_{CLC}=(\mathcal{L}_{t2i}+\mathcal{L}_{i2t})/2.\vspace{-1pt}$
Region-Text Contrastive Constraint. Inspired by the image-level contrastive learning, we attempt to construct token-wise region-text contrastive constraint using ground truth bounding box as a mask to simulate text-to-image matching. Specifically, we extract text aggregation features, i.e., the $[EOS]$ token $\bm{t}_{eos}$, from grounding encoder and compute the similarity $\bm{s}_{i}$ with each visual token $\bm{v}_{i}$ after applying normalization and an MLP projection:
(12) $\bm{s}_{i}=\sigma(<{\bm{t}_{eos}}^{\top},\ {\rm{MLP}}(\bm{v}_{i})>),\ i=1,2,...,L_{v},\vspace{-1pt}$
where $\sigma$ denotes the sigmoid function. Tokens within the bounding box are considered as positive, while those outside are regarded as negative. Subsequently, we employed Focal loss (Lin et al., 2017) and Dice/F-1 loss (Milletari et al., 2016) to constrain the aggregated similarity $\bm{s}=(\bm{s}_{1},\bm{s}_{2},...,\bm{s}_{L_{v}})$ and the nearest downsampling box mask $\bm{m}_{d}\in\mathbb{R}^{1\times H/P\times W/P}$:
(13) $\small\mathcal{L}_{RTCC}=\lambda_{focal}\mathcal{L}_{focal}(\bm{s},\bm{m}_{d})+\lambda_{dice}\mathcal{L}_{dice}(\bm{s},\bm{m}_{d}),\vspace{-1pt}$
where $\lambda_{focal}$ and $\lambda_{dice}$ are the coefficients to control the two loss functions, and $P$ is the patch size.
Training Loss. The box regression loss is formulated by leveraging smooth L1 loss (Girshick, 2015) and Giou loss (Rezatofighi et al., 2019) with coefficient $\lambda_{l_{1}}$ and $\lambda_{giou}$:
(14) $\mathcal{L}_{BOX}={\lambda_{l_{1}}}\mathcal{L}_{\text{smooth-l1 }}\big{(}\hat{\mathcal{B}},\mathcal{B}\big{)}+{\lambda_{giou}}\mathcal{L}_{\text{giou}}\big{(}\hat{\mathcal{B}},\mathcal{B}\big{)},$
where $\mathcal{B}$ donates the ground truth box. Finally, the overall training loss of the model is determined by the sum of the regression loss and the two framework constraints:
(15) $\mathcal{L}_{total}=\mathcal{L}_{BOX}+\mathcal{L}_{CLC}+\mathcal{L}_{RTCC}.$
4. Experiments
[…]
4.1. Implementation Details
Datasets and Evaluation Metrics. The effectiveness of our method is validated on five widely utilized datasets, namely the three REC datasets (RefCOCO/+/g (Yu et al., 2016; Mao et al., 2016)), as well as two Phrase Grounding (PG) datasets (ReferItGame (Kazemzadeh et al., 2014a) and Flickr30k Entities (Plummer et al., 2015)). In PG, the query pertains to a specific phrase, while in REC, the query refers to a reference expression. The text of RefCOCO+/g exhibits greater length and complexity in comparison to that of RefCOCO. We follow the previous research that employs Intersection-over-Union (IoU) as the evaluation metric. Specifically, a prediction is deemed accurate only when its IoU exceeds or equals 0.5. Finally, we compute the prediction accuracy for each dataset as a performance indicator. The detailed description and statistics information regarding these five datasets are provided in the Appendix.
Network Architecture. We employed CLIP ViT-B/16 and CLIP ViT-L/14 as the backbone of our HiVG-B and HiVG-L versions. In the based version, the Hi LoRA module utilizes a rank of 32 and an $\alpha$ coefficient of 16. The encoder layers are evenly divided into 3 groups, and Hi LoRA is applied with 3 stages accordingly. HiVG extracted 1th, 4th, 8th, and 12th layer features of the visual encoder, the cross-modal bridge injected 4th, 8th, and 12th layer, and text aggregated from 1th to 12th layer features of the text encoder. In the grounding encoder, we adopted the pre-norm instead of the post-norm structure and set the hidden dimensions as the same with text encoder. The detailed structural parameters of the framework and the large version are summarized in the Appendix.
[…]
Model update/all update Flops train test testA testA
[…]
Training Details. To prevent catastrophic forgetting, we freeze the original parameters of CLIP’s visual and textual encoders. Since the parameters of the low-stage Hi LoRA are included in the high-stage Hi LoRA, our updated parameters do not show any increase compared to the vanilla LoRA. Besides, Hi LoRA represents a PEFT approach for the pre-trained model, and the grounding encoder employs random Xavier initialization. Thus, to enhance training stability, we perform training in two stages. In the first stage, we trained the grounding encoder, regression head at a high learning rate without activating Hi LoRA. It is imperative to employ Hi LoRA for the text encoder with only one layer group as well, in order to mitigate the risk of catastrophic forgetting. The batch size is set to 80. Our model is optimized end-to-end by using the AdamW optimizer and a cosine learning scheduler with an initial learning rate of $2.5\times 10^{-4}$ for 50 epochs during the first stage. During Hi LoRA adaptation, the learning rates in three stages are $1.0\times 10^{-4}$, $0.5\times 10^{-4}$, and $0.25\times 10^{-4}$ with 20 epochs, respectively. Our framework and experiments are based on PyTorch by using 8 NVIDIA A100 GPUs. More details are provided in the Appendix.
4.2. Comparison with State-of-the-Art Methods
Experimental Setting. It is worth emphasizing that, as described in Sec. 2.1, our focus is on the transfer learning of self-supervised pre-trained models for grounding tasks. We follow the basic fine-tuning setting with the same as CLIP-VG (Xiao et al., 2023) and Dynamic-MDETR (Shi et al., 2022), etc.. In particular, we also contrast with the traditional setting of fine-tuning with pre-trained detection and language models (e.g., TransVG (Deng et al., 2021), TransVG++(Deng et al., 2023), etc.). Additionally, we also follow the previous works that utilized a mixed dataset pre-training setting (e.g., MDETR (Kamath et al., 2021), OFA(Wang et al., 2022b)) and mix the training data (only includes the RefCOCO/+/g, ReferIt, Flickr30k datasets) for intermediate pre-training. This allows us to compare our results with these works in a relatively fair manner. The details are presented in Tab. 1.
[…]
Analysis of Referring Expression Length. As shown in Fig. 4-(c), we conducted a comparison of different expression lengths on the RefCOCOg dataset. It shows that HiVG exhibits superior comprehension for longer and more complex texts, while its performance remains stable as text length increases. Furthermore, compared to CLIP-VG, our method demonstrates significantly better results.
Table 3. Ablation study of the main modules, includes Multi-layer Adaptive Cross-modal Bridge (MACB) and Hi LoRA.
[…]
val test
73.48 73.01
76.53 75.77
76.41 76.12
78.29 78.79
[…]
Architecture val test
MACB w/o. sample-agnostic weights 75.43 74.87
MACB w/o. cross-attention module 74.29 74.18
MACB w. weights’ shape $1\times 1\times H_{l}$ 74.81 74.38
MACB w. weights’ shape $n\times 1\times H_{l}$ 77.08 77.42
MACB w. weights’ shape $n\times L_{l}\times 1$ 77.42 78.49
MACB w. weights’ shape $n\times L_{l}\times H_{l}$ 78.29 78.79
MACB w. layer-to-layer linear connect 76.51 76.30
MACB w. only last layer of text features 77.07 76.82
Table 5. Ablation study of different components in Hi LoRA on RefCOCOg-test. $r$ represents the value of low rank.
[…]
Hi LoRA three-stage-$1^{th}$ ($r$=32) 76.39
Hi LoRA three-stage-$2^{th}$ ($r$=32) 77.87
Hi LoRA three-stage-$3^{th}$ ($r$=32) 78.79
Hi LoRA two-stage ($r$=32) 77.97
Hi LoRA four-stage ($r$=32) 78.16
Hi LoRA three-stage ($r$=16) 77.57
Hi LoRA three-stage ($r$=64) 76.90
Hi LoRA deep-to-shallow layer 73.93
4.3. Ablation Study
[…]
Ablation Study of MACB. As shown in Tab. 4, we conducted an ablation study on the implementation of the multi-layer adaptive cross-modal bridge (MACB, default using 12 layers of text features). The weights in the table denotes the sample-agnostic weights. The table shows that our designed structure can effectively utilize multi-level text features and achieve hierarchical adaptation. More detailed results and descriptions will be provided in the Appendix.
Ablation Study of Hi LoRA. As presented in Tab. 5 and Fig. 4-(d), we conducted an ablation study on Hi LoRA with different LoRA stages and various low ranks. It is observed that employing 3-stage Hi LoRA with low rank as 32 achieves the best performance. More detailed results will be provided in the Appendix.
[…]
Architecture of BeiT-3 (384$\times$384) val test
two-tower w. frozen 38.03 38.38
two-tower w. vanilla LoRA 64.36 64.73
two-tower w. Hi LoRA 67.11 67.39
one-tower w. frozen 44.83 46.03
one-tower w. vanilla LoRA 74.65 75.32
one-tower w. Hi LoRA 77.19 78.30
Table 7. Generality study of Hi LoRA in object detection (b) and instance segmentation (m) on MS-COCO dataset.
[…]
vanilla CLIP w. backbone frozen 640 33.4 52.6 27.7 48.2
CLIP w. vanilla LoRA 640 44.1 63.2 38.5 60.7
CLIP w. Hi LoRA 640 47.5 67.1 43.1 64.6
MaskCLIP (Dong et al., 2023) (baseline, CVPR’23) 1333 45.4 64.7 40.9 62.1
4.4. Discussion on the Generality of Hi LoRA
[…]
Figure 5. Qualitative results of our HiVG and CLIP-VG models on RefCOCOg-val datasets. We present the prediction box with IoU (in cyan) and the ground truth box (in green) in a unified image to visually display the grounding accuracy.
4.5. Qualitative Results
We visually present the results of several relatively challenging examples in Fig. 5. The attentions show the [REG] token over vision tokens from the last grounding block of each model. HiVG demonstrates exceptional semantic understanding capabilities in the complex sentences. More examples are provided in Appendix.
5. Conclusion and Future Work
In this paper, we introduce a hierarchical multimodal fine-grained modulation framework, namely HiVG, which effectively implements fine-grained adaptation of the pre-trained model in the complex grounding task. It is a concise and efficient end-to-end framework that can simultaneously alleviate two kinds of task gaps, i.e., data bias and learning objectives, through a multi-layer adaptive cross-modal bridge and a hierarchical low-rank adaptation paradigm. Our exploration in hierarchical cross-modal features offer new insights for the future grounding research, which has been neglected in past works.
In the future, as a task-agnostic hierarchical adaptation paradigm, Hi LoRA can be further investigated across diverse downstream transfer scenarios. In this paper, we only explore the implementation of a simple progressive version. Additionally, there should be further research on the settings of layer groups and LoRA stages, such as exploring the adaptive selection of the both. Finally, it is also important to explore the broader application of hierarchical LoRA for visual, linguistic, and cross-modal tasks beyond grounding and detection tasks.
Appendix
Table 8. The detailed statistics of RefCOCO (Yu et al., 2016), RefCOCO+ (Yu et al., 2016), RefCOCOg (Mao et al., 2016), ReferItGame (Kazemzadeh et al., 2014b) and Flickr30K Entities (Plummer et al., 2015) datasets. We represent test split and testA split in the same column.
Dataset Images Instances total train val test(A) testB
queries queries queries queries queries
RefCOCO (Yu et al., 2016) 19,994 50,000 142,210 120,624 10,834 5,657 5,095
RefCOCO+(Yu et al., 2016) 19,992 49,856 141,564 120,191 10,768 5,726 4,889
RefCOCOg (Mao et al., 2016) 25,799 49,822 95,010 80,512 4,896 9,602
ReferItGame(Kazemzadeh et al., 2014b) 20,000 19,987 120,072 54,127 5,842 60,103
Flickr30k (Plummer et al., 2015) 31,783 427,000 456,107 427,193 14,433 14,481
Appendix A Analysis of the Datasets
We present the statistical analysis of the five datasets employed in our experimental study. Tab. 8 presents the detailed statistics.
RefCOCO/RefCOCO+/RefCOCOg. These three datasets belong to the Referring Expression Comprehension (REC), and the images of these three datasets derived from MSCOCO (Lin et al., 2014). Expressions in RefCOCO (Yu et al., 2016) and RefCOCO+ (Yu et al., 2016) are also collected by the two-player game proposed in ReferitGame (Kazemzadeh et al., 2014b). There are two test splits called “testA” and “testB”. Images in “testA” only contain multiple people annotation. In contrast, images in “testB” contain all other objects. Expressions in RefCOCOg (Mao et al., 2016) are collected on Amazon Mechanical Turk in a non-interactive setting. Thus, the expressions in RefCOCOg are longer and more complex. RefCOCOg has “google” and “umd” splits. The “google” split does not have a public test set, and there is an overlap between the training and validation image sets. The “umd” split does not have this overlap. Therefore, we followed the previous studies (Su et al., 2023; Yu et al., 2018) and tested the RefCOCOg dataset only on the “umd” split.
ReferItGame. ReferItGame (Kazemzadeh et al., 2014b) contains images from SAIAPR12 (Escalante et al., 2010) and collects expressions through a two-player game. In this game, the first player is shown an image with an object annotation and is asked to write a natural language expression referring to the object. The second player is then shown the same image along with the written expression and is asked to click on the corresponding area of the object. If the clicking is correct, both players receive points and swap roles. If not, a new image will be presented.
Flickr30k Entities. Flickr30k Entities (Flickr30k for short)  (Plummer et al., 2015) contains images in Flickr30k dataset. The query sentences are short noun phrases in the captions of the image. The queries are simpler and easier to understand compared to RefCOCO/+/g. Therefore, the ambiguity of the expression is heightened simultaneously, resulting in a relative increase in noise.
Dataset Granularity Gaps between Pre-training and Downstream Grounding. CLIP utilizes the LAION-400M dataset (Schuhmann et al., 2021) for self-supervised pre-training, which is a noisy web dataset containing 400 million image-text pairs. As shown in Fig. 6, we present an illustration of task granularity gaps between pre-training task and downstream grounding task. It can be observed that the self-supervised pre-training typically learns coarse-grained visual and linguistic concepts from noisy web data (Fig. 6-(a)), while visual grounding requires more refined and complex interaction and alignment between linguistic and visual information (Fig. 6-(b)). The samples are derived from LAION-400M (Schuhmann et al., 2021) and RefCOCOg (Mao et al., 2016) datasets, respectively.
Figure 6. An illustration of dataset granularity gaps between pre-training task and downstream grounding task. The samples are derived from LAION-400M (Schuhmann et al., 2021) and RefCOCOg (Mao et al., 2016) datasets, respectively.
Table 9. Network structure of the proposed HiVG. params. denote the number of parameters.
Model Backbone Hidden Input Visual encoder Text encoder Grounding encoder All Update
dimension resolution layers width heads layers width heads layers width heads params. params.
HiVG-B (default) CLIP ViT-B/16 512 224 12 768 12 12 512 8 6 512 8 206M 41M
HiVG-L CLIP ViT-L/14 768 224 24 1024 16 12 768 12 6 768 8 468M 52M
Appendix B Implementation Details
Network Architecture. The detailed network structure of HiVG is shown in Tab. 9. We employ CLIP ViT-B/16 and CLIP ViT-L/14 as the backbone of our HiVG-B (default version) and HiVG-L, respectively. In the structure of HiVG-B, the image encoder and text encoder are 12-layer Transformers, while the cross-modal grounding encoder is 6-layer Transformers with the hidden embedding dimension of 512. In the structure of HiVG-L, the image encoder and text encoder are 24- and 12-layer Transformers, respectively, while the cross-modal grounding encoder is 6-layer Transformers with the hidden embedding dimension of 768. Besides, in the large version, the encoder layers are evenly divided into 4 groups, and Hi LoRA is applied with 4 stages accordingly. HiVG extracts the 6th, 12th, 18th, and 24th layer features of the visual encoder, and the cross-modal bridge is injected to the 6th, 12th, 18th, and 24th layer. We show a detailed illustration of the multi-layer adaptive cross-modal bridge (MACB) in Fig. 7.
[…]
To ensure the efficacy of contrastive learning and enhance sample diversity within a batch, we employ data shuffling to randomize the order of samples across the five datasets. The temperature coefficient $\tau$ in the contrastive learning constraint is obtained from the vanilla CLIP model. We do not use horizontal flip augmentation as it has been observed to have a detrimental impact on grounding task. Besides, other data augmentation techniques (Liao et al., 2020; Yang et al., 2020, 2019a; Deng et al., 2021; Li and Sigal, 2021) remain consistent with previous approaches.
Inference Details. Unlike previous methods, such as TransVG++, QRNet, etc., which heavily rely on high-resolution images like 640$\times$640, we adopt smaller resolution of 224$\times$224 as in the original CLIP model. To ensure compatibility, we employ a long edge alignment and short edge pad filling scheme to the image. The patch size utilized in HiVG-B and HiVG-L are 16$\times$16 and 14$\times$14. We include $[SOS]$ and $[EOS]$ token at the beginning and the end of the input text, and align it to a fixed length of 77 by padding empty tokens.
Figure 7. The detailed illustration of the multi-layer adaptive cross-modal bridge (MACB).
Model Hyperparameters. We summarize and report the hyperparameter settings of the HiVG framework in Tab. 10.
Appendix C Extra Result Analysis
Details of Figure 4-(a) of the Main Text. Inference speed (FPS) in Figure 4-(a) of the main text is measured by forwarding 5657 image-text pairs (batch size 1) from the RefCOCO testA data split through the grounding model. As many of the algorithms are no longer reproducible due to changes in the running environment, the figure is plotted based on the result in the YORO framework (Ho et al., 2023). More specifically, the FPS measurement results except for our HiVG, CLIP-VG (Xiao et al., 2023), TransVG++ (Deng et al., 2023), VG-LAW (Su et al., 2023), RCCF (Liao et al., 2020), MattNet (Yu et al., 2018) and DGA (Yang et al., 2019b), are derived from YORO (Ho et al., 2023), which by using a single Titan Xp GPU and Intel Xeon E5-2630 v4 CPU@2.20GHZ. Following YORO, the FPS of RCCF (Liao et al., 2020), MattNet (Yu et al., 2018) and DGA (Yang et al., 2019b) are copied from RCCF work (Liao et al., 2020), which measures the speed on a Titan Xp GPU (identical to YORO) and Intel Xeon E5-2680v4 CPU@2.4GHZ. For a fair comparison, we normalize the results of TransVG++ (Deng et al., 2023), VG-LAW (Su et al., 2023), CLIP-VG (Xiao et al., 2023), and our HiVG by dividing them with a factor of 3.4 to account for the higher computational capabilities of our NVIDIA A100 GPU and Intel Xeon Gold 6240R CPU@2.4GHZ setup. This normalization factor is derived from comparing the FPS achieved by TransVG on our device (i.e., 59.55, as in Table 2 of the main text) with the reported FPS in YORO (i.e., 17.51). As can be seen in the figure, both our HiVG and CLIP-VG are based on small-resolution images and achieve significantly faster inference speed. Meanwhile, our HiVG achieves the best trade-off between performance and speed.
Table 10. Hyperparameters of the HiVG framework during training. lr denotes the learning rate.
Item Value
Base model Large model
optimizer AdamW
Epoch for grounding encoder etc. 50
lr for grounding encoder etc. 2.5$\times$$10^{-4}$
weight decay 1.0$\times$$10^{-4}$
$\lambda_{l_{1}},\ \lambda_{giou}$ 2, 2
$\lambda_{focal},\ \lambda_{dice}$ 20, 2
batch size 80 32
patch size 16$\times$16 14$\times$14
low rank in Hi LoRA 32
$\alpha$ in Hi LoRA 16
Epoch for Hi LoRA 20/stage 15/stage
lr for Hi LoRA stage 1 1.0$\times$$10^{-4}$ 1.0$\times$$10^{-4}$
lr for Hi LoRA stage 2 0.5$\times$$10^{-4}$ 0.75$\times$$10^{-4}$
lr for Hi LoRA stage 3 0.25$\times$$10^{-4}$ 0.5$\times$$10^{-4}$
lr for Hi LoRA stage 4 0.25$\times$$10^{-4}$
Figure 8. Additional qualitative results of our HiVG framework on the RefCOCOg-val split. The CLIP-VG model is compared. We present the prediction box with IoU (in cyan) and the ground truth box (in green) in a unified image to visually display the grounding accuracy. We show the [REG] token’s attention over vision tokens from the last grounding block of each framework. The examples exhibit the relatively more challenging instances for grounding, thereby showcasing HiVG’s robust semantic comprehension capabilities.
Analysis of the Computational Complexity in Figure 4-(b) of the Main Text. According to Table 2 of the main text, the number of parameters in existing models (except for QRNet (Ye et al., 2022a)) is not significantly different, roughly ranging from 150M to 210M. However, the computational complexity of the Transformer architecture heavily depends on the length of input token sequences, i.e., there is an $\mathcal{O}(n^{2})$ complexity. For example, TransVG++ (Deng et al., 2023) utilizes a 640$\times$640 resolution image as input with a patch size of 16$\times$16, resulting in a sequence length of $(640/16)^{2}=40^{2}=1600$ in the vision backbone. In contrast, our HiVG employs a smaller resolution image of 224$\times$224 with a patch size of 16$\times$16; thus, our visual sequence length is only $(224/16)^{2}=14^{2}=196$. Taking [CLS] and [REG] tokens into account, HiVG’s vision sequence length is merely $(196+1)/(1600+2)=12.29\%$ compared to that of TransVG++ (i.e., TransVG++ is 8.13×8.13$\times$8.13 × larger than HiVG), demonstrating a dominant difference. Unlike the other works (Deng et al., 2023; Su et al., 2023), our framework can obtain state-of-the-art results without relying on high-resolution images. This significantly reduces the calculation complexity and greatly accelerates the training and reasoning computation of our HiVG framework.
Details of Figure 4-(d) of the Main Text. In the Figure 4-(d) of the main text, the legend for “original CLIP” represents that we only utilize the image and text encoder from vanilla CLIP as the backbone of our grounding framework while without using Hi LoRA, cross-modal bridge, and RTCC constraint etc.. Besides, it only uses the final layer of visual and text features for the grounding encoder. The legend for “CLIP w. vanilla LoRA” represents that we additionally utilize the vanilla LoRA when compare to the legend for “original CLIP”. The legend for “HiVG w/o Hi LoRA &\&& MACB” represents that our HiVG framework does not utilize the main module of Hi LoRA and MACB but utilize the RTCC constraint and multi-level visual features. The legend for “HiVG w/o Hi LoRA” represents that our HiVG framework without utilizing Hi LoRA but utilizing MACB, RTCC constraint and multi-level visual features. The legend for “HiVG w. vanilla LoRA” represents that our HiVG framework uses vanilla LoRA along with MACB, RTCC constraint and multi-level visual features. The legend for “HiVG w. Hi LoRA stage 1, 2, 3” represents our full model under the three stages of Hi LoRA.
Table 11. Ablation study of the training loss, includes Contrastive Learning Constraint (CLC) and Region-Text Contrastive Constraint (RTCC).
[…]
val test
training unstable
training unstable
77.21 77.48
78.29 78.79
[…]
val test
MACB w. sample-aware weights (with 12 layers) 77.07 77.98
MACB w. only last layer of text features 76.84 77.02
MACB w. (6th, 12th) layer of text features 77.08 77.82
MACB w. (1th, 4th, 8th, 12th) layer of text features 77.65 78.45
MACB w. (1th - 12th) layer of text features 78.29 78.79
[…]
val test
HiVG w. (12th) layer 68.69 67.43
HiVG w. (2th, 5th, 9th) layer 71.02 71.98
HiVG w. (2th, 5th, 9th, 12th) layer 71.63 72.01
HiVG w. (1th, 4th, 8th, 12th) layer 72.37 72.15
HiVG w. (3th, 6th, 9th, 12th) layer 72.08 72.04
HiVG w. (6th - 12th) layer 71.49 71.75
HiVG w. (1th - 12th) layer 71.25 71.15
Appendix D Extra Ablation Study
Ablation Study of Training Loss. As presented in Tab. 11, we extend the Table 3 of the main text, which serves as our ablation study for the two framework constraints. After the training of Hi LoRA, we observed that without the contrastive learning constraint, the performance sometimes starts to degrade or even catastrophically forgets after reaching a certain level of training accuracy. It can be seen from the table that CLC enhances stability during Hi LoRA training. Additionally, since RTCC is a token-wise constraint on the aggregated multi-level visual features, it enables a more fine-grained perception of these features.
More Detailed Ablation Results on MACB. In Table 4 of the main text, the weights in the table denotes the sample-agnostic weights. In the line 7 of Table 4 of the main text, “layer-to-layer linear connect” represents direct connect the corresponding layer of the image and text encoder by a MLP and a cross-attention module. In the line 8 of Table 4 of the main text, “only last layer of text features” represents only utilizing the last layer of text features with our multi-layer adaptive cross-modal bridge, and the shape of the sample-agnostic weights are $1\times L_{l}\times H_{l}$. As shown in Tab. 12, we provide more ablation study on the implementation of the multi-layer adaptive cross-modal bridge. In the line 1 of Tab. 12, “sample-aware weights” represents that we replace the sample-agnostic weights with a MLP structure, while also utilizing the 1th-12th layer of text features. The table shows that our designed structure can effectively select the multi-level text features and can achieve the best performance when utilizing all the 12 layer of text feature.
Ablation Study of Multi-level Visual Features. We perform an ablation study on the utilization of multi-level visual features. We conduct the ablation study on the HiVG model without utilizing all the MACB, Hi LoRA, and RTCC methods. As observed from Tab. 13, any approach that incorporates the intermediate layer of visual features outperforms solely relying on the final layer features. This confirms that some lower-level useful visual information may be discarded in the final layer, which is crucial for grounding tasks. It demonstrates that employing features from layers 1, 4, 8, and 12 yields the most favorable results.
Appendix E Additional Qualitative Results
As shown in Fig. 8, we present the grounding qualitative results with several additional challenging examples. All these results demonstrate the strong capability of our HiVG model in complex text understanding and cross-modal grounding.